OpenStudy (anonymous):

Solve by Factoring: 4x2 - 16 = 0 What are the solutions for x?

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

\[x^2 = 4\]

OpenStudy (anonymous):

\[x =\pm2\]

OpenStudy (anonymous):

or you can factor as \[4(x^2-4)=4(x+2)(x-2)\] you get the same solutions

OpenStudy (anonymous):

You technically did it more correct than me, but my first step is to always see if you can solve for x first.

OpenStudy (anonymous):

same exact thing. i only added because question said "solve by factoring"